# Anomalous diffusion from hydrodynamic recoupling in particle-hole symmetric fluids

### v1

@@ Abstract
The model is a controlled deformation of a multicomponent one-dimensional ideal gas. Particles move ballistically between collisions and stochastic symmetry-breaking events. At each collision the relative velocity is stochastically redrawn while preserving ideal-gas static correlations. Number conservation is broken by allowing momentum-conserving particle splitting and merging events at rate $\gamma_{e}$ (we reuse $\gamma_{e}$ since energy density in the Dirac fluid plays the role of particle density in the stochastic gas), while momentum conservation is broken by allowing stochastic sign flips of the center-of-mass velocity at collisions, at rate $\gamma_{p}$. In all moves, the total charge is conserved; however, at collisions the charge labels of the participating particles are randomly redrawn, subject to charge conservation, with probability $p_{\rm mix}$. This process generates a tunable Fickian contribution to charge transport. These deformations can be implemented so that the stationary state remains the grand-canonical (multicomponent) ideal gas, which we set to total particle density $\rho=1$ (with equal species split) and temperature $k_{B}T=1$, while the long-wavelength dynamics is governed by the same linear hydrodynamic equations as Eq. (3). The symmetry-breaking rates $\gamma_{p},\gamma_{e}$ are related to the microscopic probabilities $(p_{\rm flp},p_{\rm merge})$ by $O(1)$ nonuniversal coefficients which we determine empirically. Full details of the microscopic rules and the empirical determination of the symmetry breaking rates are given in the supplementary material suppmat.
@@
This behavior can be understood as an intermediate-time regime. At early times, the momentum mode is described by the linearized hydrodynamics and appears diffusive. The convective contribution to the charge transfer then scales superdiffusively with $\Delta Q(t)^{2}\sim X(t)\sim t^{3/4}$ (i.e., from repeated returns of backscattered sound waves). At longer times, however, momentum fluctuations cross over to KPZ scaling governed by the coupled Burgers equations in Eq. (Anomalous diffusion from hydrodynamic recoupling in particle-hole symmetric fluids). In this regime, both momentum and charge transfer variances are expected to scale as $t^{2/3}$. One therefore expects intermediate-time effective exponents between $1/2$ and $2/3$ for momentum transfer and between $3/4$ and $2/3$ for charge transfer, consistent with the exponents observed numerically in Fig. 4 (left). Figure 4 (right) shows the corresponding current–current correlator, which decays approximately as $t^{-1/2}$, consistent with nonintegrable long-time tails and superdiffusive charge transport.
@@
Supplementary Materials: Anomalous diffusion from hydrodynamic recoupling in particle-hole symmetric fluids
@@ I Charge transfer variance and the diffusion constant
Using charge conservation, $n(k,\omega)=\frac{k}{\omega}j(k,\omega)$, the diffusive pole (S1) implies $\mathrm{Im}\,G^{R}_{jj}(k,\omega)=\chi D\,\omega+O(\omega k^{2})$ as $k,\omega\to 0$. It follows immediately that Kubo; forster2018hydrodynamic
@@
so that the diffusion constant entering the pole of $G^{R}_{nn}$ is precisely the diffusion constant defined by the Kubo formula Kubo.
@@
Evaluating the remaining integral yields the universal asymptotic form DerridaConjecture; Roche_2005; Bertini_2015; PhysRevLett.131.210402
@@ II Singular charge diffusion from linear hydrodynamics with symmetry breaking
We focus on the convective contribution to equilibrium charge transport, which arises from the nonlinear coupling between charge and the energy current. At charge neutrality this provides a distinct contribution to equilibrium charge-transfer fluctuations, in addition to ordinary Fickian diffusion. In the weak symmetry-breaking regime, this contribution may be isolated by a decoupling assumption as described in the main text: to leading order in the symmetry-breaking rates, the sound modes evolve independently and act to advect the entire charge profile by a stochastic displacement pnas.2403327121.
@@
where $c$ is the sound velocity, $\gamma_{e}$ and $\gamma_{p}$ are the symmetry-breaking relaxation rates, and $D_{p}$ is the Fickian diffusion constant for the energy current. The noise correlators are fixed by the fluctuation-dissipation theorem landau1987fluid, and are given by
@@
$D(\gamma_{e},\gamma_{p})=\int\frac{dk}{2\pi}\,S_{pp}(0,k)=\int\frac{dk}{2\pi}\,\frac{2\gamma_{e}\bigl(c^{4}\chi_{e}k^{2}+\chi_{p}\gamma_{e}(\gamma_{p}+D_{p}k^{2})\bigr)}{\bigl[\gamma_{e}(\gamma_{p}+D_{p}k^{2})+c^{2}k^{2}\bigr]^{2}}.$ (S15)
@@
$S_{nn}(k,t)\equiv\langle n_{k}(t)n_{-k}(0)\rangle=\big\langle e^{-ikX(t)}\big\rangle.$ (S23)
@@
$S_{nn}(k,t)=\exp\!\left[-\frac{k^{2}}{2}\,\langle X(t)^{2}\rangle\right].$ (S24)
@@
$S_{nn}(k,t)\simeq\exp\!\left[-\frac{\alpha^{2}\chi_{p}}{c\sqrt{\pi}}\sqrt{\frac{t}{\gamma_{p}}}\;k^{2}\right].$ (S26)
Equivalently, $n(x,t)$ is subdiffusive with dynamical exponent $z=4$. This structure factor is characteristic of the “XNOR” or “tracer diffusion” universality class discussed in Refs. PhysRevLett.127.230602; 10.1073/pnas.2202823119; PhysRevB.106.094303, that we derived here from a purely hydrodynamic perspective.
@@ III Exact quasiparticle representation of linear fluctuating hydrodynamics
We show here that the toy model of reflected sound waves used in the main text is in fact an exact kinetic (Boltzmann) representation of the linear fluctuating hydrodynamics Eqs. (S7)–(S8). In the absence of the viscous $D_{p}\partial_{x}^{2}p$ term, the linearized equations reduce to a telegrapher-type dynamics for chiral sound packets Goldstein1951; Kac1974ASM. The sound packets can be represented in terms of quasiparticles which carry both a propagation direction and a particle/hole character, and the stochastic quasiparticle dynamics reproduces the hydrodynamic equations without approximation.
@@
We note that the kinetic description in Eq. (S28) is not unique BENZI1992145: different microscopic implementations of the Poisson processes (e.g. redistributing relaxation between the two flip channels or between explicit flips and quasiparticle birth–death noise) lead to the same closed linear equations for $e$ and $p$ with identical transport coefficients and noise correlators. This “gauge freedom” of the Boltzmann representation does not affect any hydrodynamic observables.
@@ IV Stochastic symmetry–broken ideal gas
The outgoing velocities are then $v_{1,2}^{\prime}=V\pm u^{\prime}/2$. This construction is closely analogous to the momentum–conserving Lowe–Andersen thermostat KoopmanLowe2006, in which a stochastic bath interaction thermalizes the relative motion at temperature $T$ while preserving the pair center–of–mass velocity (and hence total momentum). Unlike Lowe–Andersen, which applies such updates to nearby pairs (within a cutoff distance) at a Poisson rate independent of their relative speed, here the update is applied at actual collision events whose rate is proportional to the relative speed $|u|=|v_{1}-v_{2}|$, compensates exactly for the velocity bias in collision rates.

### v2

@@ Abstract
The model is a controlled deformation of a multicomponent one-dimensional ideal gas. Particles move ballistically between collisions and stochastic symmetry-breaking events. At each collision the relative velocity is stochastically redrawn while preserving ideal-gas static correlations. Number conservation is broken by allowing momentum-conserving particle splitting and merging events at rate $\gamma_{e}$ (we reuse $\gamma_{e}$ since energy density in the Dirac fluid plays the role of particle density in the stochastic gas), while momentum conservation is broken by allowing stochastic sign flips of the center-of-mass velocity at collisions, at rate $\gamma_{p}$. In all moves, the total charge is conserved; however, at collisions the charge labels of the participating particles are randomly redrawn, subject to charge conservation, with probability $p_{\rm mix}$. This process generates a tunable Fickian contribution to charge transport. These deformations can be implemented so that the stationary state remains the grand-canonical (multicomponent) ideal gas, which we set to total particle density $\rho=1$ (with equal species split) and temperature $k_{B}T=1$, while the long-wavelength dynamics is governed by the same linear hydrodynamic equations as Eq. (3). The symmetry-breaking rates $\gamma_{p},\gamma_{e}$ are related to the microscopic probabilities $(p_{\rm flp},p_{\rm merge})$ by $O(1)$ nonuniversal coefficients which we determine empirically. Full details of the microscopic rules and the empirical determination of the symmetry breaking rates are given in the supplementary material 2.
@@
This behavior can be understood as an intermediate-time regime. At early times, the momentum mode is described by the linearized hydrodynamics and appears diffusive. The convective contribution to the charge transfer then scales superdiffusively with $\Delta Q(t)^{2}\sim X(t)\sim t^{3/4}$ (i.e., from repeated returns of backscattered sound waves). At longer times, however, momentum fluctuations cross over to KPZ scaling governed by the coupled Burgers equations in Eq. (Noncommuting zero-noise and zero-frequency limits in particle-hole symmetric fluids). In this regime, both momentum and charge transfer variances are expected to scale as $t^{2/3}$. One therefore expects intermediate-time effective exponents between $1/2$ and $2/3$ for momentum transfer and between $3/4$ and $2/3$ for charge transfer, consistent with the exponents observed numerically in Fig. 5 (left). Figure 5 (right) shows the corresponding current–current correlator, which decays approximately as $t^{-1/2}$, consistent with nonintegrable long-time tails and superdiffusive charge transport.
@@
Supplementary Materials: Noncommuting zero-noise and zero-frequency limits in particle-hole symmetric fluids
@@ I Charge transfer variance and the diffusion constant
Using charge conservation, $n(k,\omega)=\frac{k}{\omega}j(k,\omega)$, the diffusive pole (S1) implies $\mathrm{Im}\,G^{R}_{jj}(k,\omega)=\chi D\,\omega+O(\omega k^{2})$ as $k,\omega\to 0$. It follows immediately that ,Ryogo (1957); Forster (2018)
@@
so that the diffusion constant entering the pole of $G^{R}_{nn}$ is precisely the diffusion constant defined by the Kubo formula ,Ryogo (1957).
@@
Evaluating the remaining integral yields the universal asymptotic form Derrida et al. (2004); Roche et al. (2005); Bertini et al. (2015); McCulloch et al. (2023)
@@ II Singular charge diffusion from linear hydrodynamics with symmetry breaking
We focus on the convective contribution to equilibrium charge transport, which arises from the nonlinear coupling between charge and the energy current. At charge neutrality this provides a distinct contribution to equilibrium charge-transfer fluctuations, in addition to ordinary Fickian diffusion. In the weak symmetry-breaking regime, this contribution may be isolated by a decoupling assumption as described in the main text: to leading order in the symmetry-breaking rates, the sound modes evolve independently and act to advect the entire charge profile by a stochastic displacement Gopalakrishnan et al. (2024a).
@@
where $c$ is the sound velocity, $\gamma_{e}$ and $\gamma_{p}$ are the symmetry-breaking relaxation rates, and $D_{p}$ is the Fickian diffusion constant for the energy current. The noise correlators are fixed by the fluctuation-dissipation theorem Landau and Lifshitz (1987), and are given by
@@
$D(\gamma_{e},\gamma_{p})=\int\frac{dk}{2\pi}\,S_{pp}(k,0)=\int\frac{dk}{2\pi}\,\frac{2\gamma_{e}\bigl(c^{4}\chi_{e}k^{2}+\chi_{p}\gamma_{e}(\gamma_{p}+D_{p}k^{2})\bigr)}{\bigl[\gamma_{e}(\gamma_{p}+D_{p}k^{2})+c^{2}k^{2}\bigr]^{2}}.$ (S15)
@@
$S_{nn}(k,t)\equiv\langle n_{k}(t)n_{-k}(0)\rangle=\chi\big\langle e^{-ikX(t)}\big\rangle.$ (S23)
@@
$S_{nn}(k,t)=\chi\exp\!\left[-\frac{k^{2}}{2}\,\langle X(t)^{2}\rangle\right].$ (S24)
@@
$S_{nn}(k,t)\simeq\chi\exp\!\left[-\frac{\alpha^{2}\chi_{p}}{c\sqrt{\pi}}\sqrt{\frac{t}{\gamma_{p}}}\;k^{2}\right].$ (S26)
Equivalently, $n(x,t)$ is subdiffusive with dynamical exponent $z=4$. This structure factor is characteristic of the “XNOR” or “tracer diffusion” universality class discussed in Refs. Singh et al. (2021); Nardis et al. (2022); Feldmeier et al. (2022), that we derived here from a purely hydrodynamic perspective.
Optical conductivity
Going back to the general case, the formalism outlined above can also be used to compute the optical conductivity $\sigma(\omega)$ given by the Kubo formula
$\sigma(\omega)=\lim_{k\to 0}\int_{0}^{\infty}dt{\rm e}^{i\omega t}S_{jj}(k,t),$ (S27)
with $S_{jj}(k,t)$ the current-current correlator. Using the continuity equation $\partial_{t}n+\partial_{x}j=0$, we can compute the $k\to 0$ limit of the current-current correlator using the general form
$S_{nn}(k,t)=\chi\exp\!\left[-D_{\rm Fick}k^{2}t-\frac{k^{2}}{2}\,\langle X(t)^{2}\rangle\right].$ (S28)
of the structure factor. Note that the variable $X(t)$ remains Gaussian in the general case, as it is a linear combination of Gaussian variables given by linearized hydrodynamics. We find
$\sigma(\omega)=\chi D_{\rm Fick}+\frac{\chi}{2}\int_{0}^{\infty}dt{\rm e}^{i\omega t}\frac{d^{2}}{dt^{2}}\langle X(t)^{2}\rangle,$ (S29)
where the first term corresponds to the Fickian contribution, whereas the second one is due to convective terms. Finally, using the general expression (7) and focusing on the real part of the conductivity, we obtain
$\sigma(\omega)=\chi D_{\rm Fick}+\chi\int\frac{dk}{2\pi}S_{pp}(k,\omega).$ (S30)
In the d.c. limit $\omega\to 0$, we recover $\sigma=\chi D$ with $D=D_{\rm Fick}+\int\frac{dk}{2\pi}S_{pp}(k,0)$ the diffusion constant.
@@ III Exact quasiparticle representation of linear fluctuating hydrodynamics
We show here that the toy model of reflected sound waves used in the main text is in fact an exact kinetic (Boltzmann) representation of the linear fluctuating hydrodynamics Eqs. (S7)–(S8). In the absence of the viscous $D_{p}\partial_{x}^{2}p$ term, the linearized equations reduce to a telegrapher-type dynamics for chiral sound packets Goldstein (1951); Kac (1974). The sound packets can be represented in terms of quasiparticles which carry both a propagation direction and a particle/hole character, and the stochastic quasiparticle dynamics reproduces the hydrodynamic equations without approximation.
@@
We note that the kinetic description in Eq. (S32) is not unique Benzi et al. (1992): different microscopic implementations of the Poisson processes (e.g. redistributing relaxation between the two flip channels or between explicit flips and quasiparticle birth–death noise) lead to the same closed linear equations for $e$ and $p$ with identical transport coefficients and noise correlators. This “gauge freedom” of the Boltzmann representation does not affect any hydrodynamic observables.
@@ IV Stochastic symmetry–broken ideal gas
The outgoing velocities are then $v_{1,2}^{\prime}=V\pm u^{\prime}/2$. This construction is closely analogous to the momentum–conserving Lowe–Andersen thermostat Koopman and Lowe (2006), in which a stochastic bath interaction thermalizes the relative motion at temperature $T$ while preserving the pair center–of–mass velocity (and hence total momentum). Unlike Lowe–Andersen, which applies such updates to nearby pairs (within a cutoff distance) at a Poisson rate independent of their relative speed, here the update is applied at actual collision events whose rate is proportional to the relative speed $|u|=|v_{1}-v_{2}|$, compensates exactly for the velocity bias in collision rates.